6x= -3(2x-4): 6x= -3(2x-4)
Answer by checkley77(3654)

6x=-3(2x-4)
6x=-6x+12
6x+6x=12
12x=12
x=12/12
x=1 answer.
6*1=-3(2*1-4)
6=-3(2-4)
6=-3*-2
6=6